# Permutation groups, partition lattices and block structures

### v2

@@ 4.4 The embedding theorem
The Krasner–Kaloujnine theorem [25] says that, if $G$ is a transitive but imprimitive permutation group, then $G$ is embeddable in the wreath product of two groups which can be extracted from $G$ (the stabiliser of a block acting on the block, and $G$ acting on the set of blocks).
@@
For the second part, we use the fact that, for a given point $\alpha\in\Omega$, there is a natural correspondence between partitions and certain subgroups of $G$ containing $G_{\alpha}$, where the partition $\Pi$ corresponds to the setwise stabiliser of the part of $\Pi$ containing $\alpha$; meet and join correspond to intersection and product of subgroups. Let $H_{1}$, $H_{2}$, $P$, $P^{-}$ be the subgroups corresponding to $\Phi_{1}$, $\Phi_{2}$, $\Pi$, $\Pi^{-}$. Then the definition of $\mathcal{G}(m)$ shows that $H_{i}\cap P=P^{-}$ for $=1,2$, while the partitions $\Phi_{i}\vee\Pi$ correspond to the subgroups $H_{i}P$. The actions we are interested in are thus $H_{i}P$ on the cosets of $H_{i}$. We have
@@
for $=1,2$; so coset representatives of $P^{-}$ in $P$ are also coset representatives for $H_{i}$ in $H_{i}P$. Thus we have a natural correspondence between these sets. Since $H_{1}\leq H_{2}$, we have $H_{1}P\leq H_{2}P$, and the result holds.  $\Box$
@@
The proof uses properties of distributive lattices: we deal with some of these first. Since these lemmas are not specifically about lattices of partitions, we depart from our usual convention and use lower-case italic letters for elements of a lattice, and 00 and $1$ for the least and greatest elements respectively.
@@
Suppose that $L$ is the lattice of down-sets in a poset $M$. Let $p$ be a minimal element of $M$ (so that $\{p\}$ is a down-set). Then the interval [$[\{p\},1]$ in $L$ is isomorphic to the lattice of down-sets in $M\setminus\{p\}$.
@@
Let $z=\{p\}$. Let $\mathop{\mathrm{JI}}(L)$ be the set of join-indecomposables in $L$. We construct an order-isomorphism $F$ from $\mathop{\mathrm{JI}}(L)\setminus\{z\}$ to $\mathop{\mathrm{JI}}([z,1])$.
The map $F$ is defined by
@@
for $a\in\mathop{\mathrm{JI}}(L)\setminus\{z\}$. We have to show that it is a bijecton and preserves order. First we show that its image is contained in $\mathop{\mathrm{JI}}([z,1])$.
@@
Now we turn to the proof of Theorem 4.22. The proof is by induction on the number of elements in $M$. We take $\Pi_{0}$ to be a minimal non-$E$ partition, corresponding to a minimal element $p\in M$. We decorate things computed in the interval $[\Pi_{0},U]$ with bars; for example, $\bar{G}^{*}(q)$ corresponds to the group associated in this lattice with the element $q\neq p$ (which is not in general the same as $G^{*}(q)$). Thus $\bar{G}$ is the group induced by $G$ on the set of parts of $\Pi_{0}$, which is a PB group with associated poset $M\setminus\{p\}$; our induction hypothesis will imply that the group $\bar{G}$ is embedded in the generalised wreath product of the groups $\bar{G}^{*}(q)$ for $q\in M\setminus\{p\}$.
Let $\Pi$ be a join-indecomposable partition in $[\Pi_{0},U]$, corresponding to the element $q\in M\setminus\{p\}$. As we saw in the proof of Lemma 4.24, there are two possibilities:
  • Case 1: $\Pi$ is join-indecomposable in the lattice $L$. Then $\Pi^{-}$ is above $\Pi_{0}$, and so the group $\bar{G}^{*}(q)$ is the same as $G^{*}(q)$.
  • Case 2: $\Pi=\Pi_{0}\vee\Psi$, where $\Psi$ is join-indecomposable in $L$ and $\Pi_{0}$ is not below $\Psi$. Consder the set $\bar{\mathcal{G}}(q)$, where the bar denotes that it is computed in the lattice $[\Pi_{0},U]$. A partition $\Phi$ belongs to this set if it is above $\Pi_{0}$ and satisfies $\Phi\wedge\Pi=\bar{\Pi}^{-}$, where again the bar denotes the unique maximal element below $\Pi$ in $[\Pi_{0},U]$. An easy exercise shows that $\bar{\Pi}^{-}\wedge\Psi=\Psi^{-}$; hence
@@ 4.5 Intersections of posets
Given a family $(G(i):m_{i}\in M)$ of groups indexed by a set $M$, any partial order on $M$ gives rise to a generalised wreath product of the groups. So we have a map from partial orders on $M$ to generalised wreath products of the groups $G(i)$. In this section, we prove that this map preserves order and intersections. To explain the terminology, inclusions and intersections of partial orders on the same sets are given by inclusions and intersections of the sets of ordered pairs comprising the order relations. It is easy to show that the intersection of partial orders is a partial order.
@@
Let $(G(i):m_{i}\in M)$ be a family of groups indexed by a set $M$, and let $\mathcal{M}_{1}=(M,\sqsubseteq_{1})$ and $\mathcal{M}_{2}=(M,\sqsubseteq_{2})$ be two posets based on $M$. Then
@@
A linear extension of a poset $M$ is a total order which includes the poset. It is a standard result that a poset is the intersection of all its linear extensions. (If $i$ is below $j$ in the poset, then $i$ is below $j$ in every linear extension. Conversely, if $i$ and $j$ are incomparable, then there are two linear extensions, in one of which $i$ is below $j$, and in the other $j$ is below $i$.)
@@ 5.1 Computing questions
$\begin{array}[]{|c|rrr|}\hline\cr n&\hbox{Trans}&\hbox{OB}&\hbox{PP}\\ \hline\cr 10&45&44&42\\ 11&8&8&8\\ 12&301&285&276\\ 13&9&9&9\\ 14&63&62&59\\ 15&104&104&102\\ 16&1954&1886&1833\\ 17&10&10&10\\ 18&983&922&900\\ 19&8&8&8\\ 20&1117&1100&1019\\ \hline\cr\end{array}$
@@
In fact, there is a GAP package by Bamberg, Hanaki and Lansdown which can be used to check isomorphism. Using this package, we could add a column to the above table giving the number of different association schemes which result (and identifying them in the Hanaki–Miyamoto tables).
@@ 5.2 Some problems
1. Is it true that the generalised wreath product of groups with the OB property has the OB property? (We saw in Proposition 4.7 that direct and wreath product preserve the OB property.)
@@
Acknowledgements. We are grateful to Michael Kinyon for drawing our attention to the paper [36].

### v3

@@ 4.4 The embedding theorem
The Krasner–Kaloujnine theorem [27] says that, if $G$ is a transitive but imprimitive permutation group on $\Omega$, then $G$ is embeddable in the wreath product of two groups which can be extracted from $G$ (the stabiliser of a block acting on the block, and $G$ acting on the set of blocks).
@@
For the second part, we use the fact that, for a given point $\alpha\in\Omega$, there is a natural correspondence between partitions and certain subgroups of $G$ containing $G_{\alpha}$, where the partition $\Pi$ corresponds to the setwise stabiliser of the part of $\Pi$ containing $\alpha$; meet and join correspond to intersection and product of subgroups. Let $H_{1}$, $H_{2}$, $P$, $P^{-}$ be the subgroups corresponding to $\Phi_{1}$, $\Phi_{2}$, $\Pi$, $\Pi^{-}$. Then the definition of $\mathcal{G}(m)$ shows that $H_{i}\cap P=P^{-}$ for $i=1,2$, while the partitions $\Phi_{i}\vee\Pi$ correspond to the subgroups $H_{i}P$. The actions we are interested in are thus $H_{i}P$ on the cosets of $H_{i}$. We have
@@
for $i=1,2$; so coset representatives of $P^{-}$ in $P$ are also coset representatives for $H_{i}$ in $H_{i}P$. Thus we have a natural correspondence between these sets. Since $H_{1}\leq H_{2}$, we have $H_{1}P\leq H_{2}P$, and the result holds.  $\Box$
@@
The proof uses properties of distributive lattices: we deal with some of these first. Since these lemmas are not specifically about lattices of partitions, we depart from our usual convention and use lower-case italic letters for elements of a lattice, and 0 and $1$ for the least and greatest elements respectively.
@@
Remark. The identity in Lemma 4.22 is commonly known as the cancellation property for distributive lattices, which appears in [16] as Exercise 6.6. Note that distributivity and the cancellation property are in fact equivalent lattice properties, but we only use one direction here, so we only prove the direction we use. The converse can be proved using Theorem 2.5. In particular, it is clear that $P_{5}$ and $N_{3}$ do not admit cancellation.
@@
Suppose that $L$ is the lattice of down-sets in a poset $M$. Let $p$ be a minimal element of $M$ (so that $\{p\}$ is a down-set). Then the interval $\{p\},1]$ in $L$ is isomorphic to the lattice of down-sets in $M\setminus\{p\}$.
@@
Let $z=\{p\}$. Let $\mathop{\mathrm{JI}}(L)$ be the set of join-indecomposables in $L$. Since lattices are generated by their join-indecomposable elements, it suffices to construct an order-isomorphism $F$ from $\mathop{\mathrm{JI}}(L)\setminus z$ to $\mathop{\mathrm{JI}}([z,1])$. The map $F$ is defined by
@@
for $a\in\mathop{\mathrm{JI}}(L)\setminus\{z\}$. We have to show that it is a bijection and preserves order. First we show that its image is contained in $\mathop{\mathrm{JI}}([z,1])$.
@@
Now we turn to the proof of Theorem 4.21. The proof is by induction on the number of elements in $M$. We take $\Pi_{0}$ to be a minimal non-$E$ partition, corresponding to a minimal element $p\in M$. We decorate things computed in the interval $\Pi_{0},U]$ with bars; for example, $\bar{G}^{*}(q)$ corresponds to the group associated in this lattice with the element $q\neq p$ (which is not in general the same as $G^{*}(q)$). Thus $\bar{G}$ is the group induced by $G$ on the set of parts of $\Pi_{0}$, which is a PB group with associated poset $M\setminus\{p\}$; our induction hypothesis will imply that the group $\bar{G}$ is embedded in the generalised wreath product of the groups $\bar{G}^{*}(q)$ for $q\in M\setminus\{p\}$.
Let $\Pi$ be a join-indecomposable partition in $\Pi_{0},U]$, corresponding to the element $q\in M\setminus\{p\}$. As we saw in the proof of Lemma 4.23, there are two possibilities:
  • Case 1: $\Pi$ is join-indecomposable in the lattice $L$ of downsets of $M$. Then $\Pi^{-}$ is above $\Pi_{0}$, and so the group $\bar{G}^{*}(q)$ is the same as $G^{*}(q)$.
  • Case 2: $\Pi=\Pi_{0}\vee\Psi$, where $\Psi$ is join-indecomposable in $L$ and $\Pi_{0}$ is not below $\Psi$. Consider the set $\bar{\mathcal{G}}(q)$, where the bar denotes that it is computed in the lattice $\Pi_{0},U]$. A partition $\Phi$ belongs to this set if it is above $\Pi_{0}$ and satisfies $\Phi\wedge\Pi=\bar{\Pi}^{-}$, where again the bar denotes the unique maximal element below $\Pi$ in $\Pi_{0},U]$. An easy exercise shows that $\bar{\Pi}^{-}\wedge\Psi=\Psi^{-}$; hence
@@ 4.5 Intersections of posets
Given a family $(G(i)\leq\mathop{\mathrm{Sym}}(\Omega_{i})\,:\,m_{i}\in M)$ of transitive permutation groups indexed by a set $M$, any partial order on $M$ gives rise to a generalised wreath product of the groups. So we have a map from partial orders on $M$ to generalised wreath products of the groups $G(i)$. In this section, we prove that this map preserves order and intersections. To explain the terminology, inclusions and intersections of partial orders on the same sets are given by inclusions and intersections of the sets of ordered pairs comprising the order relations. It is easy to show that the intersection of partial orders is a partial order.
@@
Let $(G(i)\leq\mathop{\mathrm{Sym}}(\Omega_{i})\,:\,m_{i}\in M)$ be a family of transitive permutation groups indexed by a set $M$, and let $\mathcal{M}_{1}=(M,\sqsubseteq_{1})$ and $\mathcal{M}_{2}=(M,\sqsubseteq_{2})$ be two posets based on $M$. Then
@@
A linear extension of a poset $M$ is a total order which includes the poset. It is a standard result that a poset is the intersection of all its linear extensions. (If $i$ is below $j$ in the poset, then $i$ is below $j$ in every linear extension. Conversely, if $i$ and $j$ are incomparable, there is a linear extension in which $i$ is below $j$, and one in which $j$ is below $i$.)
@@ 5.1 Computing questions
[…]
@@
In fact, there is a GAP package [10] by Bamberg, Hanaki and Lansdown which can be used to check isomorphism. Using this package, we could add a column to the above table giving the number of different association schemes which result (and identifying them in the Hanaki–Miyamoto tables).
@@ 5.2 Some problems
1. Under what conditions is the generalised wreath product of OB groups OB?